• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Kalyan

    Log entry time 16:08:21 on July 4,2010

    Entry number 330031

    keyword=Day shift summary

    Shift summary
    -----------------
    08.00 : Controlled access to check coincidence trigger (Bogdan)
    10.05 : Access completed. Asked MCC for beam.
    10.20 : Septum turned on ( config: charlie), confirmed corrector magnet currents with MCC.
    10.40 : MCC is calibrating ion chambers for optics target.
    11.40 : Escorted access
    11.41 : Septum off
    12.20 : Beam permit. Septum on.
    12.50 : spot++ looks like y direction is off. Asked MCC to go to y= -2.0
    13.10 : moved to BPM y = +2
    14.00 : CHL tripped. no beam for at least 4 hours.


    Run target current comments
    -------------------------------------
    1662 optics 0.6uA Raster check ( only T3)
    1663 optics 0.6uA only T3
    1664 optics 0.6uA Raster check (only T1)
    1665 optics 0.6uA only T3
    1666 optics 5uA T3 (sieve in)
    1671 optics 5uA T3 "
    1672 optics 5uA T1 "
    1673 optics 5uA T1 "
    1674 optics 5uA T1 "
    1675 optics 5uA T1
    1676 optics 5uA T1 " , BPM Y =-2